OpenStudy (anonymous):

Use the Rational Root Theorem to list all possible rational roots of the polynomial equation x3 - 4x2 + 7x - 8 = 0. Do not find the actual roots. A. –8, –1, 1, 8 B. –8, –4, –2, –1, 1, 2, 4, 8 C. 1, 2, 4, 8 D. no possible roots

OpenStudy (anonymous):

i think its D not sure tho

OpenStudy (anonymous):

possible rational zeros of g(x)=x3-4x2+x-10 using the rational zero theorem? ... [ possible rational roots of the following equation:3x^4 + 5x^3 + 3x^2 - 7x + 9 = 0

OpenStudy (anonymous):

so would the answer be B

OpenStudy (anonymous):

ya

OpenStudy (tamtoan):

can 1 be the root of that equation ? :)

OpenStudy (anonymous):

yes

OpenStudy (tamtoan):

if you sub x = 1 into the equation : x^3 - 4x^2 + 7x - 8 = 0, do you have both side equal ?

OpenStudy (anonymous):

i dont think so

OpenStudy (tamtoan):

isn't A, B, C all have 1 in it ? :)

OpenStudy (anonymous):

yes